OpenStudy (anonymous):

what is the derivative of 2x(x-2)^3?

OpenStudy (anonymous):

if you don't want to expand (which is not that hard) and use the power rule, you can use the product rule

OpenStudy (anonymous):

How would I use the power rule? Do i distribute the 2x?

OpenStudy (anonymous):

expanding gives \(2 x^4-12 x^3+24 x^2-16 x\) and that is easy to find the derivative of

OpenStudy (anonymous):

but you can use \[(fg)'=f'g+g'f\] with \[f(x)=2x,f'(x)=2,g(x)=(x-2)^3,g'(x)=3(x-2)^2\]

OpenStudy (anonymous):

Would you mind walking me through that? I'm not quite sure how you got the end product.

OpenStudy (anonymous):

you mean how i got this\[2x(x-2)^3=2 x^4-12 x^3+24 x^2-16 x\]??

OpenStudy (anonymous):

Yea, that would be awesome.

OpenStudy (anonymous):

\[2x(x-2)^3=2x(x-2)(x-2)(x-2)\] is a start then either grind it til you find it, or recall that \((a-b)^3=a^3-b^3=a^3-3 a^2 b+3 a b^2-b^3\)

OpenStudy (anonymous):

typo there, i meant \[(a-b)^3=a^3-3 a^2 b+3 a b^2-b^3\]

OpenStudy (anonymous):

so you get \[(x-2)^3=x^3+6x^3+12x-8\]

OpenStudy (anonymous):

then multiply everything by \(2x\) and get the answer i wrote above

OpenStudy (anonymous):

Oh, I see. Thank you so much. Breaking down the problem makes it much simpler.

OpenStudy (anonymous):

yw btw you can still use the product rule, then you don't have to do all that multiplication

OpenStudy (anonymous):

\[f(x)=2x(x-2)^3= 2x^4-12x^3+24x^2-16x f'(x)=8x^3-36x^2+48x-16\]